OpenStudy (anonymous):

4x+6y=8 GIVE ANY NUMBER FOR x and y. ( from -10 to 10) other than x=2 and y=0

OpenStudy (gw2011):

x=-10 and y=8 Proof: (4)(-10)+6(8)=-40+48=8